OpenStudy (anonymous):

A car covers 72 kilometers in the first hour of its journey. In the next hour, it covers 90 kilometers. What is the amount of work done by the car? The total mass of the car, including its passengers, is 2.5 × 103 kilograms

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hello, my name is Stephen and my major is physics. 2.8 × 105 joules is the amount of work done by the car. Solution: Work = Force times displacement = change in kinetic energy but first convert the 90 km/h and 72 km/h into m/s (multiply each by 1000 and divide by 3600 - you'll find it to be 25 [ m/s and 20 m/s respectively) change in KE = (1/2)(2.5x10^3)(25^2) - (1/2)(2.5x10^3)(20^2) = 281,250 J which can be written in scientific notation as 2.8 x 10^5 joule